OpenStudy (anonymous):

Is the answer *none*??? Q)For what value of a is following equation true? lim x->inf ((x+a)/(x-a)) = e

OpenStudy (shubhamsrg):

unless a=f(x) .. i dont see any value of a to be satisfying that..

OpenStudy (anonymous):

bcuz someone ACTUALLY plugged in (1+(1/x))^x for e value.... i mean, for another question

OpenStudy (anonymous):

i looked up similar question online, but it wasn't ((x+a)/(x-a)) but ((x+a)/(x-a))^x so when that person plugged in the e value, both sides (^x) cancalled out

OpenStudy (anonymous):

it's not *none* right? bcuz that'd b too easy to be true.

OpenStudy (shubhamsrg):

i see,,if there is ^x over there,,ans according to me becomes a=1/2 ..

OpenStudy (anonymous):

yea! i was wondering if my professor made a typo. cuz this question is way too easy.... a has to be none no matter wat...

OpenStudy (sirm3d):

\[\large a=1/2\]

OpenStudy (anonymous):

for lim x-> inf ((x+a)/(x-a)) = e ?????????

OpenStudy (sirm3d):

if there's no ^x, the limit is 1, not e.

OpenStudy (anonymous):

but the question asks: For what value of a is following equation true

OpenStudy (anonymous):

which means without ^x, the limit is 1. which is not 0. which means answer is none right?

OpenStudy (sirm3d):

yes.

OpenStudy (anonymous):

phewww cool!

OpenStudy (sirm3d):

now, if ^x is supposed to be there, the answer to the problem is a=1/2

OpenStudy (anonymous):

yay good thing that i had this prob haha